OpenStudy (anonymous):

find the exact value of cos(u-v) given that sin u=-9/41 and cos v=15/17. (Both u and v are in quadrant IV.)

OpenStudy (anonymous):

cos(u-v)=cosu cosv + sinu sinv............ sin v nd cos u are not known ... find by cos(360-x)=sinx nd sin(360-x)=-cosx